A wheel accelerates with a constant angular acceleration of 6.5 rad s-2


. If the initial angular velocity is 2 rad s-1

.


What is the angular velocity at t = 2 s.


1
Expert's answer
2021-10-10T15:55:59-0400
"\omega_f=2+6.5(2)=15\frac{rad}{s}"